a solution is prepared from 2.4gram of a organic compound and 75 gram of benzene. if the freezing point of the solution is -0.975 centigrade.what will be the molecular formula of the organic compound if the organic compound is made of carbon and hydrogen.
1
Expert's answer
2013-07-09T09:51:58-0400
Theformula for freezing point depression (for non-electrolytes like sucrose) is
ΔTf = Kf * m, where ΔTf = change in freezing point, Kf = molal freezing pointdepression constant for the solvent, and m = molality of the solution.
freezing point of benzene is 5.5 oC
ΔTf = 5.5 + 0.975 = 6.475
The Kf for benzene is 5.12 degrees Celsius C/m.
m = ΔTf / Kf;
m = 6.475 / 5.12 = 1.26
The molality of the solution:
molality = moles solute / mass (kg) solvent
moles solute = molality * mass (kg) solvent
moles solute = 1.26 * 0.075 kg = 0.095 moles
molar mass = mass / moles
molar mass = 2.4g / 0.095 moles = 25.26 g/mol
molecular formula of the organic compound is C2H2